OpenStudy (anonymous):

1. For each of your three trials state the following: • heat needed to melt the ice (q) • enthalpy of fusion • percent error from the accepted enthalpy of fusion of water of 334 J/g

OpenStudy (anonymous):

79.72 cal/gm is the Heat of fusion

OpenStudy (anonymous):

) To heat 1 kg (about 1 liter) of water from 283.15 K to 303.15 K (10 °C to 30 °C) requires 83.6 kJ. However, to melt ice also requires energy. To heat ice from 273.15 K to water at 293.15 K (0 °C to 20 °C) requires: (1) 333.55 J/g (heat of fusion of ice) = 333.55 kJ/kg = 333.55 kJ for 1 kg of ice to melt PLUS (2) 4.18 J/(g·K) × 20K = 4.18 kJ/(kg·K) × 20K = 83.6 kJ for 1 kg of water to increase in temperature by 20 K = 417.15 kJ

OpenStudy (anonymous):

@magicalsd

OpenStudy (anonymous):

you probably needed some information